OpenStudy (anonymous):

How would I write an equation of a line in standard form with a slope of -8 and passing through the point (-10, 2).

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

put it in the point-slope form, since you are given a point and a slope

OpenStudy (anonymous):

hey Misty:] so what would that look like?

OpenStudy (misty1212):

do you know what the point slope form looks like?

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (misty1212):

yeah plug in the numbers in the the appropriate spots

OpenStudy (anonymous):

it says in standard form so i'm confused?

OpenStudy (misty1212):

you are not there yet you have to do this first

OpenStudy (misty1212):

go ahead plug the numbers in to the point slope form, then we can write it in "standard form"

OpenStudy (anonymous):

y- 2=-8(x- -10)

OpenStudy (misty1212):

ok good now distribute on the right

OpenStudy (anonymous):

im lost I don't know how I would get this in standard form?

OpenStudy (misty1212):

we are working towards that be patient dear it takes a few steps \(\color\magenta\heartsuit\)

OpenStudy (misty1212):

use the distributive property to remove the parentheses on the right side of the equal sign

OpenStudy (anonymous):

y- 2=-8x- (-10)

OpenStudy (misty1212):

also you probably want to start with \[y-2=-8(x+10)\]

OpenStudy (misty1212):

now distribute the \(-8\) on the right side

OpenStudy (misty1212):

you know what i am saying? \[a(b+c)=ab+ac\] is the distributive law

OpenStudy (misty1212):

so what is \[-8(x+10)\]?

OpenStudy (anonymous):

-8x+10

OpenStudy (misty1212):

oh no lets try again

OpenStudy (misty1212):

\[a(b+c)=ab+ac\] not \(ab+c\)

OpenStudy (anonymous):

-80x +10

OpenStudy (misty1212):

you just erased the parentheses, you did not distribute

OpenStudy (misty1212):

\[-5(x+3)=-5x-15\] for example

OpenStudy (anonymous):

-8x -80

OpenStudy (misty1212):

yay

OpenStudy (misty1212):

\[y-2=-8x-80\] now add \(2\) to both sides

OpenStudy (anonymous):

y=-8x+82

OpenStudy (misty1212):

hmm what is \(-80+2\)?

OpenStudy (anonymous):

-78

OpenStudy (misty1212):

ok so one more step we have \[y=-8x-72\]

OpenStudy (misty1212):

to put in in standard form, add \(8x\) to both sides

OpenStudy (anonymous):

8x+y=-78

OpenStudy (misty1212):

now we are done!

OpenStudy (anonymous):

that's it? haha you made it so easy!:]

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

thank you so much hun!:]